Find 6°48'59? + 28°19'36?. Write your answer in simplest terms.
Answer by nerdybill(5413)

6°48'59" + 28°19'36"
.
Adding the "seconds":
59+36 = 95 seconds
Or, 1 min 35 seconds
.
Adding the "minutes" (including the 1 from above)
1 + 48 + 19 = 68 minutes
Or, 1 deg 8 minutes
.
Adding the "degrees" (including the 1 from above)
1 + 6 + 28 = 35 degrees
.
Pulling it all together:
35°8'35"